# https://kartik-3004.github.io/facexformer_web/

### v1

@@ Appendix 0.A Overview
As part of the supplementary material, we present the following as an extension to the ones shown in the paper:
  • Cross-dataset Evaluation (Section 6)
  • Parameter and FPS Analysis (Section 0.C)
  • In-the-wild Visualization (Section 0.D)
  • Additional Results (Section 0.E)
  • Additional Training details (Section 0.F)
Appendix 0.B Cross-Dataset Evaluation
Method 300VW LFWA FFHQ(Age) FFHQ(Gender)
NME Acc. Acc. Acc.
HyperFace [51] 8.50 - - 91.90
AllinOne [52] 7.28 - 34.88 80.83
FaceXFormer 9.83 72.23 46.60 94.92
Table 6: Cross Dataset evaluation of FaceXFormer in comparison to existing multi-task models
@@ Appendix 0.B Cross-Dataset Evaluation
We conduct additional cross-dataset experiments to demonstrate the effectiveness of FaceXFormer in scenarios closely resembling real life, where face images are previously unseen, unconstrained, and characterized by significant variability in background, lighting, pose, and other factors. From Table 6, it is evident that FaceXFormer surpasses existing multi-task baselines in classification tasks. However, its effectiveness in landmark detection under in-the-wild scenarios is only moderate. The paper provides an explanation for this behavior, which we reiterate here: HyperFace [51] and AllInOne [52] are primarily trained on regression-based tasks, which boosts their performance on individual regression tasks. Consequently, they might outperform FaceXFormer in landmark detection but fall short in other tasks that only FaceXFormer can handle. The cross-dataset results echoes our rationale presented in the paper: the need for a unified facial analysis model capable of performing multiple tasks on in-the-wild, unconstrained faces, particularly for real-life applications. FaceXFormer addresses this gap and demonstrates competitive performance across various tasks.
Appendix 0.C Parameter and FPS Analysis
Model Params GFLOPs FPS B4FPS
Swin-B 91.98 17.98 39.89 150.25
Swin-B (seg) 91.06 17.93 41.09 156.19
Swin-B (reg) 91.01 17.54 41.54 164.77
Swin-B (cls) 91.47 17.56 41.18 162.02
SwinV2-B 92.14 19.60 30.91 117.69
SwinV2-S 54.08 12.19 31.32 118.77
SwinV2-T 32.70 7.42 54.58 198.20
ConvNext-B 92.80 17.90 73.09 185.69
ConvNext-S 54.57 11.17 74.75 246.71
ConvNext-T 32.93 6.94 105.02 336.47
ResNet101 48.23 10.54 66.84 238.02
Table 7: Parameter and FPS analysis with different backbones for FaceXformer.
The parameter and FPS analysis for different backbones in the FaceXformer framework provides insightful trends regarding computational efficiency and performance, as shown in Table 7. Notably, for tasks such as regression (reg), classification (cls), and segmentation (seg), there is a minimal change in parameters and frames per second (FPS), which underscores the efficiency of the pipeline, as conventional face segmentation pipelines generally have more parameters and lower FPS than classification and regression pipelines. For example, Swin-B for segmentation, regression, and classification tasks demonstrates remarkable consistency in performance with FPS rates of 41.09, 41.54, and 41.18, respectively, while maintaining similar parameter counts around 91 million. In terms of overall FPS, ConvNext backbones outperform the others, with the ConvNext-T model achieving an 105.02 FPS, followed by ResNet101 and then Swin models.
Appendix 0.D In-the-wild Visualization
We randomly select images from the publicly available FFHQ [29] dataset and treat them as "in-the-wild" images, which were not used for training. We input the image along with multiple task tokens and present the qualitative results for all the tokens in Figure 5. We observe that FaceXFormer delivers competitive results across a variety of faces with variations in age, gender, race, head pose, background, lighting, etc.
@@ Appendix 0.D In-the-wild Visualization
Figure 5: Visualization of "in-the-wild" images queried for multiple task tokens. Attributes represent the $40$ binary attributes defined in the CelebA [42] dataset, indicating the presence ($1$) or absence (00) of specific facial attributes.
Appendix 0.E Additional results
In this section, we expand upon the results presented in the main paper by providing class-wise F1 scores for face parsing on the LaPa [41] and CelebAMask-HQ [34] datasets. Additionally, we elaborate on the results of landmark detection across various subsets of these datasets, utilizing multiple metrics, as detailed in Table 10.
FaceXFormer achieves the highest F1-score for the "skin" class and the lowest for the "left-eyebrow" on the LaPa dataset. Similarly, on CelebAMask-HQ, it records the highest F1-score for "skin" and the lowest for the "right-eyebrow." From Tables 8 and 9, it is apparent that the class-wise segmentation performance ranking of FaceXFormer aligns with the pattern observed in other baseline models. Furthermore, FaceXFormer notably surpasses the recent state-of-the-art model, FP-LIIF [56], on the LaPa dataset, even with a lower input resolution.
From Table 10, we observe that FaceXFormer surpasses existing multi-task baseline models, particularly HyperFace and AllInOne, with a NME of $3.52$ and $4.81$ on the 300W [53] and WFLW [76] datasets, respectively. It also outperforms FaRL [88] on the 300W challenge set, achieving an NME of $5.13$ compared to FaRL’s $5.19$. Additionally, FaceXFormer exceeds FaRL’s performance on the WFLW dataset with a $\text{FR}_{@10\%}$ of 4.56, compared to $5.72$ for FaRL. Furthermore, its performance is on par with the current state-of-the-art model, demonstrating its capacity to generalize across various heterogeneous tasks without sacrificing performance in any specific task.
Method Input Res. Skin Hair Nose LE RE LB RB LL IM UL Mean
EAGR [65] 473 97.3 96.2 97.1 89.5 90.0 86.5 87.0 89.0 90.0 88.1 91.1
AGRNET [64] 473 97.7 96.5 97.3 91.6 91.1 89.9 90.0 90.1 90.7 88.5 92.3
DML-CSR [87] 473 97.6 96.4 97.3 91.8 91.5 90.4 90.4 89.9 90.5 88.0 92.4
FaRL [88] 512 97.1 93.0 97.2 91.6 91.5 90.1 89.6 89.1 89.4 87.2 91.6
FP-LIIF [56] 256 97.1 95.8 97.0 85.5 86.2 88.8 88.6 87.8 89.4 85.0 90.1
FaceXFormer 224 96.7 92.7 96.1 91.5 91.4 88.0 88.1 88.5 89.8 86.6 91.1
Table 8: Class-wise segmentation results on LaPa dataset
Method Input Res. Skin Hair Nose LE RE LB RB LL IM UL Mean
EAGR [65] 473 96.2 94.9 94.0 88.6 88.7 85.7 85.2 91.2 95.0 88.9 90.84
AGRNET [64] 473 96.5 95.2 93.9 88.7 89.1 85.5 85.6 91.1 92.0 89.1 90.67
DML-CSR [87] 473 95.7 94.5 93.9 89.4 89.6 85.5 85.7 91.0 91.8 87.4 90.45
FaRL [88] 512 96.1 94.9 93.8 89.0 89.0 85.3 85.4 90.0 91.7 88.1 90.34
FP-LIIF [56] 256 96.4 95.1 93.7 88.5 88.5 84.5 84.3 90.3 92.1 87.5 90.09
FaceXFormer 224 96.1 94.4 93.4 89.2 89.0 84.4 84.3 90.0 89.3 87.2 89.83
Table 9: Class-wise segmentation results on CelebAMask-HQ dataset
Method 300W WFLW
Full Common Challenge Full Pose Exp Ill Makeup Occl Blur
NME NME NME NME $\text{FR}_{@10\%}$ $\text{AUC}_{@10\%}$ NME NME NME NME NME NME
Wing [17] - - - 4.99 7.56 0.532 8.75 5.36 4.93 5.41 6.37 5.81
SBR [14] 4.10 3.28 7.58 - - - - - - - - -
DeCaFa [12] 3.39 2.93 5.26 4.62 4.84 0.563 8.11 4.65 4.41 4.63 5.74 5.38
HRNet [69] 3.32 2.87 5.15 4.60 4.64 - 7.94 4.85 4.55 4.29 5.44 5.42
STAR Loss [90] 2.87 2.52 4.32 4.02 2.32 0.605 6.76 4.27 3.97 3.84 4.80 4.58
FaRL [88] 3.35 2.90 5.19 4.80 5.72 0.545 8.78 5.09 4.74 4.99 6.01 5.35
HyperFace [51] 3.62 3.22 3.62 5.12 6.09 0.509 9.54 5.82 5.41 5.31 6.43 6.14
AllinOne [52] 4.52 3.98 6.67 6.90 15.4 0.391 13.4 7.91 7.05 7.42 8.53 7.95
FaceXFormer 3.52 3.13 5.13 4.81 4.56 0.530 8.80 5.67 5.12 5.07 6.00 5.86
Table 10: Landmarks detection results on various subsets of 300W and WFLW
Appendix 0.F Datasets and Implementation Details
@@
0.F.1 Face Parsing
We use LaPa [41] and CelebAMask-HQ [34] for training and evaluation of FaceXFormer. LaPa consists of around 22,000 facial images with 11-category pixel level map. CelebAMask-HQ contains 30,000 high-resolution face images annotated with 19 classes. To use both datasets for training, we take the intersection of class labels from both datasets. The classes used for training and evaluation include: skin, face, nose, left eye, right eye, left eyebrow, right eyebrow, upper lip, mouth, and lower lip. While training, we perform augmentations with random rotation (±$\pm 18^{\circ}$), random scaling (±$\pm 10\%$), random translation ($1\%\times 224$), random horizontal flip ($50\%$), random gray ($10\%$), random Gaussian blur ($10\%$) and random gamma adjustment($10\%$). We perform Tanh warping [40] on the LaPa dataset to ensure that the network can segment the whole face image while focusing on the face region. On CelebAMask-HQ, instead of Tanh-warping, we perform a resizing operation to $224\times 224$, since all face images are already aligned.
0.F.2 Landmarks Detection
We utilize the 300W [53] and WFLW [76] datasets for the training and evaluation of FaceXFormer. The 300W and WFLW datasets contain 3,148 and 7,500 images in their respective training sets. The 300W dataset includes 689 test images, categorized into three overlapping test sets: common (554 images), challenge (135 images), and full (689 images). It includes a large variety of identities, expression, illumination conditions, pose, occlusion, and face size. The images are annotated with 68 landmark points. The WFLW dataset includes 2,500 images for testing, each annotated with 98 landmark points. For the sake of unification, we convert the 98 landmarks to 68 landmarks using the standard landmark mapping. Additionally, it includes annotations for occlusion, head pose, makeup, illumination, blur, and expressions. We employ the 300VW [57] dataset for cross-dataset testing of multi-task methods. It offers three test sets: Category-A (well-lit conditions, comprising 31 videos with 62,135 frames), Category-B (mildly unconstrained conditions, with 19 videos containing 32,805 frames), and Category-C (challenging conditions, including 14 videos with 26,338 frames). We report the mean of these three categories. During training we apply augmentations such as random rotation (±$\pm 18^{\circ}$), random scaling (±$\pm 10\%$), random translation ($5\%\times 224$), random horizontal flip ($50\%$), random gray ($20\%$), random Gaussian blur ($30\%$), random occlusion ($40\%$) and random gamma adjustment($20\%$). Additionally, we align the images using five landmarks points.
0.F.3 Head Pose Estimation
We utilize the 300W-LP [93] and AFLW2000 [94] datasets, which contain approximately 122,000 samples and 2,000 images, respectively. We evaluate performance using the BIWI [16] dataset, comprising 15,678 images of 20 individuals (6 females and 14 males, with 4 individuals recorded twice). The head pose range covers about ±$\pm 75^{\circ}$ yaw and ±$\pm 60^{\circ}$ pitch. During training, we loosely crop the face images based on the landmarks and apply augmentations, including random gray ($10\%$), random Gaussian blur ($10\%$), and random gamma adjustment($10\%$).
0.F.4 Attributes Recognition
@@
0.F.5 Age/Gender/Race Estimation
We employ the FairFace [28] and UTKFace [85] datasets for training, and the FFHQ [29] dataset for testing. FairFace comprises 108,501 images, balanced across seven racial groups: White, Black, Indian, East Asian, Southeast Asian, Middle Eastern, and Latino. The UTKFace dataset contains 20,000 facial images annotated with age, gender, and race. In our work, we follow the ’race-4’ annotation scheme, categorizing individuals into five racial labels: White, Black, Indian, Asian, and Others. Age annotations are categorized into decade bins: 0-9, 10-19, 20-29, 30-39, 40-49, 50-59, 60-69, and over 70. Gender is annotated with two labels: male and female. During training, we apply augmentations, such as random rotation (±$\pm 18^{\circ}$), random scaling (±$\pm 10\%$), random translation ($1\%\times 224$), random horizontal flip ($50\%$), random gray ($10\%$), random Gaussian blur ($10\%$), and random gamma adjustment($20\%$).
0.F.6 Landmarks Visibility Prediction
We utilize the COFW [5] dataset, which is annotated with 29 landmarks for landmarks visibility prediction. Each landmark is associated with 29 binary labels that indicate its visibility. We loosely crop the faces and apply augmentations, including random horizontal flip ($50\%$), random gray ($10\%$), random Gaussian blur ($10\%$), and random gamma adjustment($20\%$).

### v2

@@ Appendix A Overview
As part of the Appendix, we present the following as an extension to the ones shown in the paper:
  • Cross-dataset Evaluation (Section B)
  • Ablation study (Section C)
  • In-the-wild Visualization (Section D)
  • Dataset details (Section E)
Appendix B Cross-Dataset Evaluation
@@
We conduct additional cross-dataset experiments to demonstrate the effectiveness of FaceXFormer in scenarios that closely resemble real-life conditions. These scenarios involve previously unseen, unconstrained face images characterized by significant variability in background, lighting, pose, and other factors. As shown in Table B.1, FaceXFormer outperforms the existing state-of-the-art model, STARLoss [125], on the 300VW dataset. This highlights FaceXFormer’s effectiveness in landmark detection under in-the-wild scenarios. The cross-dataset results support the rationale presented in this paper: the necessity of a unified facial analysis model capable of performing multiple tasks on unconstrained, in-the-wild faces, particularly for real-time applications. FaceXFormer addresses this gap and achieves state-of-the-art performance.
Method 300VW (Cat.A) 300VW (Cat.B) 300VW (Cat.C) LFWA (Gender)
NME NME NME Acc.
PANDA [118] - - - 92.00
STARLoss [125] 3.97 3.39 8.42 -
FaceXFormer 3.90 3.58 6.75 92.74
Table B.1: Cross Dataset evaluation of FaceXFormer.
Appendix C Ablation Study
To evaluate the contribution of each component in FaceXFormer, we conduct an ablation study focusing on the importance of specific design choices and their impact on performance across various tasks. Specifically, we perform experiments by: a) Excluding multi-scale features and relying solely on the final-layer features. b) Removing the FXDec decoder from the architecture. The results of these experiments are summarized in Table C.1.
Method HPE Lnd Attr. Age
F1 MAE Acc. MAE
w/o multi-scale 3.70 4.70 91.21 4.21
w/o FXDec 17.16 31.49 79.90 16.15
FaceXFormer 3.52 4.67 91.83 4.17
Table C.1: Impact of various components of FaceXFormer on performance.
From the results, it is evident that the FXDec decoder, which incorporates self-attention and interactions between face tokens and task tokens, plays a critical role in performance. Without FXDec, there is a significant drop across all tasks. For instance, the Mean Absolute Error (MAE) for landmark detection increases from $4.67$ to $31.49$, and the accuracy for attribute classification drops from $91.83$% to $79.90$%. This substantial decline highlights the decoder’s importance in effectively capturing complex feature relationships necessary for these tasks. Similarly, excluding multi-scale features leads to reduced performance across all tasks, with a particularly notable impact on head pose estimation and age estimation. The F1 score for head pose estimation increases from $3.52$ to $3.70$ (indicating worse performance since lower is better), and the MAE for age estimation increases from $4.17$ to $4.21$. These results shows the importance of integrating multi-scale features to capture both global and local information essential for accurate predictions.
Appendix D In-the-wild Visualization
We randomly selected images from the publicly available CelebAMask-HQ [38] dataset and treated them as “in-the-wild” images for tasks it was not specifically trained on. We present the qualitative results for all the tokens in Figure D.1. Our observations indicate that FaceXFormer produces promising results. However, we noted inconsistencies in age estimation and race prediction.
@@ Appendix D In-the-wild Visualization
Figure D.1: Visualization of ”in-the-wild” images for multiple tasks. Attributes represent the $40$ binary attributes defined in the CelebA [50] dataset, indicating the presence ($1$) or absence (00) of specific facial attributes.
Appendix E Datasets and Implementation Details
@@
E.1 Face Parsing
We use CelebAMask-HQ [38] for training and evaluation of FaceXFormer. CelebAMask-HQ contains 30,000 high-resolution face images annotated with 19 classes. The classes used for training and evaluation include: skin, face, nose, left eye, right eye, left eyebrow, right eyebrow, upper lip, mouth, and lower lip. During training, we resize the images to $224\times 224$, before feeding them into the model.
E.2 Landmarks Detection
We utilize the 300W dataset [75] for the training and evaluation of FaceXFormer. The 300W dataset contains 3,148 images in its training set and 689 test images, which are categorized into three overlapping test sets: common (554 images), challenge (135 images), and full (689 images). It encompasses a wide variety of identities, expressions, illumination conditions, poses, occlusions, and face sizes. All images are annotated with 68 landmark points. For cross-dataset testing of multi-task methods, we employ the 300VW dataset [78]. This dataset provides three test categories: Category-A (well-lit conditions, comprising 31 videos with 62,135 frames), Category-B (mildly unconstrained conditions, consisting of 19 videos with 32,805 frames), and Category-C (challenging conditions, including 14 videos with 26,338 frames). We report the results for all three categories. During training, we apply various data augmentations such as random rotation (±$\pm 18^{\circ}$), random scaling (±$\pm 10\%$), random translation ($5\%\times 224$), random horizontal flip ($50\%$), random gray ($20\%$), random Gaussian blur ($30\%$), random occlusion ($40\%$) and random gamma adjustment($20\%$). Additionally, we align the images using five landmarks points.
E.3 Head Pose Estimation
We utilize the 300W-LP dataset [128], which contains approximately 122,000 samples. For performance evaluation, we use the BIWI dataset [18], comprising 15,678 images of 20 individuals (6 females and 14 males, with 4 individuals recorded twice). The head pose range spans approximately ±$\pm 75^{\circ}$ yaw and ±$\pm 60^{\circ}$ pitch. During training, we loosely crop the face images based on the landmarks and apply several augmentations, including random gray ($10\%$), random Gaussian blur ($10\%$), random resized crop ($80\%to100\%$)and random gamma adjustment($10\%$).
E.4 Attributes Prediction
@@
E.5 Age/Gender/Race Estimation
We utilize the FairFace [32] and UTKFace [120] datasets for training, and the FFHQ [33] dataset for cross-dataset testing. FairFace comprises 108,501 images, balanced across seven racial groups: White, Black, Indian, East Asian, Southeast Asian, Middle Eastern, and Latino. The UTKFace dataset contains 20,000 facial images annotated with age, gender, and race. In our work, we follow the ’race-4’ annotation scheme, categorizing individuals into five racial labels: White, Black, Indian, Asian, and Others. Age annotations are categorized into decade bins: 0–9, 10–19, 20–29, 30–39, 40–49, 50–59, 60–69, and over 70. Gender is annotated with two labels: male and female. Additionally, we incorporate the MORPH-II dataset [72], which contains 55,134 facial images of 13,617 subjects aged between 16 and 77 years. This dataset provides annotations for age, gender, and race, with a predominance of male subjects and a significant representation of Black and White individuals. For age estimation tasks, we train on both UTKFace and MORPH-II datasets and evaluate our models on the MORPH-II dataset to assess performance. During training, we apply augmentations such as random rotation (±$\pm 18^{\circ}$), random scaling (±$\pm 10\%$), random translation ($1\%\times 224$), random horizontal flip ($50\%$), random grayscale conversion ($10\%$), random Gaussian blur ($10\%$), and random gamma adjustment ($10\%$).
E.6 Facial Expression Recognition
We utilize the RAF-DB [42] and AffectNet [58] datasets for training and RAF-DB [42] dataset for intra-dataset evaluation. RAF-DB is a facial expression dataset with approximately 30,000 images. The dataset includes variability in subjects’ age, gender, ethnicity, head poses, lighting conditions, and occlusions (e.g., glasses, facial hair, or self-occlusion). RAF-DB provides annotations for seven basic emotions that are surprise, fear, disgust, happiness, sadness, anger, and neutral. AffectNet is one of the largest facial expression datasets with approximately 440,000 images that are manually annotated for the presence of eight discrete facial expressions: neutral, happy, angry, sad, fear, surprise, disgust, contempt. During training, we apply augmentations such as random rotation (±$\pm 18^{\circ}$), random scaling (±$\pm 10\%$), random translation ($1\%\times 224$), random horizontal flip ($50\%$), random grayscale conversion ($10\%$), random Gaussian blur ($10\%$), random color jitter ($10\%$), and random gamma adjustment ($10\%$).
E.7 Visibility Prediction
We utilize the COFW [7] dataset, which is annotated with 29 landmarks for landmarks visibility prediction. Each landmark is associated with 29 binary labels that indicate its visibility. We loosely crop the faces and apply augmentations, including random horizontal flip ($50\%$), random gray ($10\%$), random Gaussian blur ($10\%$), and random gamma adjustment($10\%$).